OpenStudy (anonymous):

-3y=-x-7 graph the eqaution by plotting three points? Please help, thank you.

OpenStudy (mww):

This is a line. A line follows the same path no matter what. In fact, you can draw a line if you know just 2 points that line on that line. How do you do that? Well substitute three x values and find out what the corresponding y values are. First: -3y=-x-7 can be converted into -3y = -(x+7) so \[y=\frac{ x+7 }{ 3 }\] Now this makes it a lot easier. Let x be any 3 values. x = -1, 0 and 2 say When x = -1, y = 6/3 = 2 When x = 0, y = 7/3 When x = 2, y = 9/3 = 3 So there are our three points (-1,2), (0,7/3) and (2,3) Plot them join them and you have a line.

OpenStudy (anonymous):

no fractions are allowed

OpenStudy (mww):

Well then choose another x value that generates an integer y value. x = 5 gives y =12/3 = 4, (5,4) being the point. Technically you only need 2 points to make a line. Fractions might make it harder to plot but they still lie on that line if they satisfy the equation

OpenStudy (anonymous):

can u help meh with another one???